OpenStudy (anonymous):

Find the sum to n terms of 2+10+30+68+130.........?

OpenStudy (anonymous):

again! lol

OpenStudy (anonymous):

@satellite73 Plzz Explain wat u did yesterday

OpenStudy (anonymous):

i have No idea....lol...Plzz

OpenStudy (anonymous):

ok slowly but fast cause i have to go first i looked at the successive differences and discovered that the third differences is a constant, namely 6 that tells me that the formula for each term (not the sum) will be cubic

OpenStudy (anonymous):

Each number : \[n ^{3}+n\]

OpenStudy (anonymous):

then i tried to figure out what that would look like, knowing that it involves a cube patter in pretty clear \(2=1^3+1,10=2^2+2,30=3^3+3...\) meaning each term look like \[a_k=k^3+k\] now this i did by eyeballs, although there are methods

OpenStudy (anonymous):

then i added, knowing two formulas \[\sum_{k=1}^nk^3=\left(\frac{n(n+1)}{2}\right)^2\] and \[\sum_{k=1}^nk=\frac{n(n+1)}{2}\] and that is all

OpenStudy (anonymous):

Yup....that makes..Sense....and gives...me the Idea......Thxxx..).....Buddy

OpenStudy (anonymous):

that tells me that \[\sum_{k=1}^n(k^3+k)=\left(\frac{n(n+1)}{2}\right)^2+\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

this...question haunted me...all night..)

OpenStudy (anonymous):

yw btw there are methods to find the cubic polynomial if you know the terms and know that it is a cube frankly i don't really know them, but this one was real easy to suss out